OpenStudy (broskishelleh):

Task 4—Solving Rational Equations Using the equation below as a model, fill in numbers in the place of a and b to create a rational equation that has an extraneous solution. Part 1. Show all work to solve for x in the equation and check the solution. Part 2. Explain how to identify the extraneous solution and what it means. Part 3. Use complete sentences to explain how the remainder theorem is used to determine whether your linear binomial is a factor of your polynomial function

OpenStudy (broskishelleh):

@Nurali

OpenStudy (broskishelleh):

Help

OpenStudy (broskishelleh):

@superRbow15

OpenStudy (broskishelleh):

@ikram002p

OpenStudy (broskishelleh):

@phi

OpenStudy (broskishelleh):

Phi please help me

OpenStudy (phi):

You have to do this Using the equation below as a model, fill in numbers in the place of a and b to create a rational equation that has an extraneous solution. What equation do they show you?

OpenStudy (broskishelleh):

None, you have to make one

OpenStudy (phi):

Using the equation below as a model<--- ??

OpenStudy (broskishelleh):

Oh wait they do, give me a sec

OpenStudy (broskishelleh):

@phi

OpenStudy (phi):

they say fill in numbers in the place of a and b where is "b" ?

OpenStudy (broskishelleh):

X = b here

OpenStudy (broskishelleh):

They didn't design the course correctly and it has errors

OpenStudy (phi):

do you mean \[ \frac{x+a}{ax} = b \]?

OpenStudy (phi):

so far, you have not posted an equation with a and b in it.

OpenStudy (phi):

can you find and post the "model equation"?

OpenStudy (broskishelleh):

@phi I am so sorry my father wanted me to get him something from the store

OpenStudy (broskishelleh):

x plus a over ax = b over x \[ \frac{ x+ a }{ ax } = \frac{ b }{ x}\]

OpenStudy (broskishelleh):

DONE!

OpenStudy (broskishelleh):

Can we use this @phi

OpenStudy (phi):

yes, but they want you to replace a and b with numbers

OpenStudy (broskishelleh):

So, can I put random numbers?

OpenStudy (broskishelleh):

\frac{ x+ 1 }{ (1)x } = \frac{ 2 }{ x}

OpenStudy (broskishelleh):

\[\frac{ X + 1 }{ 1(x) } = \frac{ 2 }{ x }\]

OpenStudy (phi):

ok, a=1 and b=2 \[ \frac{ x+ 1 }{ (1)x } = \frac{ 2 }{ x} \]

OpenStudy (broskishelleh):

Yep

OpenStudy (phi):

we could cross multiply to get x(x+1) = 2x x^2 + x = 2x x^2 -x = 0 x(x-1)=0 x=0 or x=1 the x=0 is extraneous, because we are not allowed to divide by 0 (which is what happens in the original equation if x is 0)

OpenStudy (broskishelleh):

Is dis de answer?

OpenStudy (phi):

yes